OpenStudy (anonymous):

A Chemistry Technician measures 2.50 grams of sodium sulfate and adds enough water to make a final volume of 500 mL of solution. what is the molarity of this solution (M=Moles of solute/Liters of solution)

OpenStudy (anonymous):

n= m/M and then: c= n/V

OpenStudy (anonymous):

Molar Mass of Na2SO4 = 23 x 2 + 32 x 1 + 16 x 4 = 142g Moles of Na2SO4 = 2.50g Na2SO4 x ( 1mol Na2SO4 / 142gNa2SO4) Moles of Na2SO4 = 0.0176 moles of Na2SO4 Now, Convert 500ml in to litres Litre of solution = 0.5L Molarity = Mole/ litre solution = 0.15 / 0.5 = 0.035M The molarity of the solution is 0.035M